Find a minimal representation for the system G (s) s2 = [ s+1 (s+2)(8+3) $2+38+2 1
Answer #1

G(s)=[\frac{s^{2}}{s^{2}+3s+2} \ \ \ \frac{(s+1)}{(s+2)(s+3)}]

G1(s) =\frac{s^{2}}{s^{2}+3s+2}

$2 Gl(s) (s +1)(s +2)

=\frac{A}{(s+1)}+\frac{B}{(s+2)}

A=\frac{s^{2}}{s+2},\ \ \ s=-1

A=1

B=\frac{s^{2}}{s+1},\ \ \ s=-2

B=-4

=\frac{1}{(s+1)}+\frac{-4}{(s+2)}

=e^{-t}-4e^{-2t}

  

G2(s)=\frac{(s+1)}{(s+2)(s+3)}

=\frac{A}{(s+2)}+\frac{B}{(s+3)}

A=\frac{s+1}{s+3},\ \ \ s=-2

A=-1

B=\frac{s+1}{s+2},\ \ \ s=-3

B=2

=\frac{-1}{(s+2)}+\frac{2}{(s+3)}

=-e^{-2t}+2e^{-3t}

G(s)=G1(s)*G2(s)

=(e^{-t}-4e^{-2t})(-e^{-2t}+2e^{-3t})

=-e^{-t}*e^{-2t} -4e^{-2t} *e^{-2t}+ 2e^{-3t}*e^{-t} +2e^{-3t}*-4e^{-2t}

=-e^{-3t} -4e^{-4t} + 2e^{-4t} -8e^{-5t}

=-e^{-3t} -2e^{-4t} -8e^{-5t}

=-e^{-3t}(1 +2e^{-t} +8e^{-3t} )